OpenStudy (anonymous):

any help, natural logs −10ln(p−8)=−10

zepdrix (zepdrix):

Hey cola :) \[\Large\rm -10 \ln(p-8)=-10\]Dividing each side by -10 gives us:\[\Large\rm \ln(p-8)=1\]

zepdrix (zepdrix):

Here is how we rewrite logs in exponential form: